solve y'=x+y by power series method of solution
1
Expert's answer
2020-06-21T16:52:42-0400

Let, solution is

"y=\sum_{n=0}^{\infty}a_nx^n"

Thus,

"y'=\sum_{n=1}^{\infty}a_nx^{n-1}"

Hence, on simplification the differential equation

"y'=y+x"

We get,

"a_0=a_1,a_2=\frac{a_0+1}{2}, a_3=\frac{a_0+1}{6}"

and

"\frac{a_n}{a_{n-1}}=\frac{1}{n}, \forall n\geq 3"

Thus,

"a_n=\frac{6}{n!}, \forall n\geq 3"

Hence, solution will be,

"y=a_0+a_1x+(\frac{a_0+1}{2})x^2+(a_0+1)\sum_{n=3}^{\infty}\frac{1}{n!}\\
\implies y=a_0+a_1x+(\frac{a_0+1}{2})x^2+(a_0+1)(e^x-1-x-\frac{1}{2}x^2)\\
y=(a_0+1)e^x-x-1"